# New Online Communities: Graph Deep Learning on Anonymous Voting Networks to Identify Sybils in Polycentric Governance

### v4

@@ Abstract
This research examines the polycentric governance of digital assets in blockchain-based Decentralized Autonomous Organizations (DAOs). It offers a theoretical framework and addresses a critical challenge facing decentralized governance by developing a method to identify sybils, or spurious identities. The method uses graph deep learning techniques to identify sybil activity in a DAO governance dataset (snapshot.org). Specifically, a Graph Convolutional Neural Network (GCNN) learned voting behaviours and a fast k-means vector clustering algorithm (FAISS) used the high dimensional embeddings to identify similar nodes in a graph. The results reveal that deep learning can effectively identify sybils, reducing the voting graph by 2-5%. This research underscores the importance of sybil resistance in DAOs and offers a novel perspective on decentralized governance, informing future policy, regulation, and governance practices.
@@ 1   Introduction
Decentralized Autonomous Organizations (DAOs) represent a critical intersection of social science and emerging technologies. Originating from blockchain architectures, DAOs have transcended their initial role in cryptocurrency governance to become complex social systems that defy traditional classification. In an attempt to elucidate their underlying structure and functions, this research offers a theoretical framework for DAOs and presents an inductive method to identify sybils, thereby tackling a substantial issue facing the decentralized governance of digital assets.
In this research, a theoretical exploration of endogenous and exogenous models drawn from research on Online Communities, coupled with empirical data, shapes the foundation of the neoinstitutional theory offered here. In a nod to Elinor Ostrom’s work, this theory posits that DAOs can effectively govern Digital Common Pool Resources (DCPRs) through polycentric governance, despite their idiosyncratic challenges, if they effectively manage sybils. Sybils are fraudulent identities used to manipulate voting within digital decentralized governance. While traditional governance frameworks readily address the ‘one person, one vote’ principle, the anonymity provided by blockchain technologies complicates the influence of sybils in decentralized governance.
To address this governance challenge, we used graph deep learning techniques to identify sybil activity on a popular DAO governance forum, snapshot.org. Specifically, our method involves a Graph Convolutional Neural Net (GCNN) with multiple engineered feature layers that learned vector node embeddings from a voting graph to construct a “similarity” subgraph, where similar nodes are identified using a fast k-means clustering algorithm (Facebook AI Similarity Search, FAISS) and relabelled. The similarity graph is then reduced by combining predicted sybils. While this method is highly inductive and sensitive to feature development and learning hyperparameters, experimental results suggest that deep learning techniques can meaningfully detect sybils and the resulting subgraph of ‘true’ identities is approximately 2-5% reduced. Social scientists, policy makers, system designers, and security professionals should consider these results when studying, designing, and implementing online communities and ensure that vulnerable governance practices are made resilient to intentional and unintentional sybil attacks.
@@ 2   Deciphering DAOs: Advancing Research in Digital Collective Action
The key unifying factor among all DAOs is their use of, or responsibility for, a decentralized ledger, or “blockchain.” Central to each DAO is a shared digital asset on a blockchain, often referred to as a “token.” Recognizing the pivotal role of “token economics” in these novel, community-led platforms, we find a compelling area of research: the exploration of how these digitally augmented human collectives manage these shared assets. This exploration draws our attention towards the concept of a “Common Pool Resource” [37] and my suggested adaptation, the “Digital Common Pool Resource” (DCPR). The study of DCPRs, as a result, could offer valuable insights into the nature and operation of DAOs.
The importance of DAO research also goes beyond academic interest. It offers an opportunity to challenge and extend our understanding of collective action, digital resource management, and decentralized governance. Further, studying DAOs offers insights into ongoing, cutting-edge experiments attempting to create robust, long-enduring institutional forms beyond the control (and regulation) of nation states. Moreover, these efforts buttress new and existing cryptocurrencies and frustrate criminal investigations, which routinely seek cryptocurrency traces when investigating crime. More optimistically and progressively, through detailed data analysis we can better comprehend the intricacies of DAOs and tap into their potential to reshape our digital environments, perhaps even as a form of “plural public” [29].
@@ 3   Emergence and Evolution of DAO Studies
As the discourse within DAO Studies evolves, we are starting to see a shift from insular, participant-derived narratives towards the integration of established socio-economic theories that offer a broader context. A pivotal theoretical framework used to analyze these phenomena has been Hirschman’s paradigm of “Exit, Voice, and Loyalty” [28]. In this model, ‘Exit’ refers to the option of leaving an organization or community when its quality deteriorates or when disagreements arise, while ‘Voice’ involves expressing dissatisfaction or working towards change from within the organization. ‘Loyalty,’ on the other hand, reflects a commitment to the organization that might inhibit ‘Exit’ and promote ‘Voice.’
@@
Similarly, many scholars have addressed how trust is transformed by DAOs. Much literature on the role of trust in blockchain technology presupposes that its function is to lower transaction costs within organizations, a sentiment summarized by Williamson [54] with his famous theory of Transaction Cost Economics (TCE). Halaburda et al.’s [25] research extends this argument by exploring how blockchain, as a trust-enhancing technology, can alter the dynamics of transaction governance in organizations, including DAOs. This work explores the intricacies of how blockchain’s inherent features, such as decentralization, transparency, and immutability, contribute to reducing uncertainties and transaction costs that are traditionally managed through hierarchical organizational structures. Halaburda’s insights are particularly relevant in the context of DAOs, which are thought to operate on principles of decentralized governance and collective decision-making. By reducing the need for intermediaries and centralized control, blockchain technology enables DAOs to operate with a higher degree of trust among participants, fostering a collaborative environment that is less susceptible to opportunistic behaviors, a key concern in TCE.
Moreover, Halaburda’s perspective aligns with Hardin’s theory of trust, emphasizing how trust in DAOs is not merely a function of self-interest or moral commitment but also heavily influenced by the technological framework that underpins these organizations. This viewpoint is further developed by Lemieux [30], who reconstructs Hardin’s theory [27] to incorporate the role of blockchain technology in shaping trust dynamics. A different model of trust emerges from Hardin’s theory of trust, where trust is “grounded in the truster’s assessment of the intentions of the trusted with respect to some action,” which is, typically, based in self-interest, moral commitment, and other idiosyncratic factors. These motivations are clearly present in the organizational dynamics of DAOs, however without alteration Hardin’s theory fails to address the increasing role of technology. Others, such as Wright and De Filippi [55], have focused on the ways technology transforms trust, drawing a parallel to the substitutability, or lack therefore, of law by code.
@@ 3.1   Quantitative and Empirical DAO Research
We performed HHI (see Figure 1), Gini (see Figure 2) and $n$-top players (see Figure 3) analyses on our dataset of 7.2m votes.
@@
In terms of governance, in the Snapshot dataset Bellavitis, Fisch, and Momtaz found that 15% of DAOs have a token ownership threshold to make governance proposals, 160 votes on average per voter, and the proposal approval rate has hovered around 90%. However, these are extremely naive counts that do not discount the considerable noise in DAO governance nor the obvious role of sybils, so while the results offer a helpful baseline, we question the accuracy of the findings. For instance, using slightly less naive counting methods on a much larger Snapshot dataset of 59,939 proposals, we find approximately 100 average votes per voter.
@@ 3.2   Existing Research Challenges for DAO Studies
The theoretical underpinnings of DAO studies are relatively undeveloped, with the initial body of literature often drawing from a ‘native’ theory of technological utopianism that is self-referential and uncritical. This self-referential use of theory may have lead to an echo chamber effect that focused attention on decentralization and autonomy at the expense of richer social phenomena.
@@ 4   Online Communities
5   Towards Polycentric Governance
@@
5.1   Digital Common Pool Resources
Ostrom’s ideas about polycentric governance can be adopted and adapted to digital assets and their governance. Here, we present the novel idea of “Digital Common Pool Resources” (DCPRs) as a digital adaptation of her theory of Common Pool Resources (CPRs). In this context, resources are not physical but digital, like codebases, online communities, or cryptographic assets. Ostrom’s principles find renewed relevance in the rapidly evolving world of blockchain and cryptocurrency, where decentralized, community-based control is paramount. Scholars such as De Filippi and Hassan [22] have further explored how blockchain technologies align with Ostrom’s principles, offering new opportunities for cooperative governance. Ostrom’s neoinstitutional approach provides a foundational framework for understanding self-governance in decentralized digital systems, focusing on the human-centered aspects of these technologies. This approach emphasizes collective action, reciprocity, and shared governance, aligning closely with the core philosophies of decentralization in the digital age.
@@ 5.1   Digital Common Pool Resources
By applying these principles to the decentralized governance of Web3, the creation and management of digital commons can reflect the successful self-governance of tangible resources. Our approach, described below, helps mitigate the security challenges of digital spaces, ensuring that these commons are managed effectively and sustainably.
@@ 6.1   Ethical considerations
It should be noted that using deep learning methods to analyze anonymous users on public financial and data records poses significant ethical risks and such methods should be considered a form of social network deanonymization. Even though the data are ostensibly public, individuals may not have consented or even be aware that their financial transactions could be linked back to them in an identifiable way. This becomes particularly troubling given that financial data can reveal sensitive aspects of an individual’s life—spending habits, income sources, and social networks—that they may not want publicly disclosed. In light of previous findings [16] on ethical research risks for cryptocurrencies, even when lacking real “human subjects,” research without consent exacerbates ethical concerns. Moreover, researchers undertaking this type of analysis must be acutely aware that their methods could be weaponized to infringe upon privacy or even perpetrate crimes.
@@ 6.2   Sybil Resistance and Defence in Depth for Community Design
These efforts, largely emerging from computer science, parallel the evolution of Ostrom-style commons management practices emerging in DAOs. DAOs continue to experiment with alternative identity systems, including Proof of Participation (PoP), Decentralized IDs (DIDs), and Self Sovereign IDs (SSIDs). New identity systems that guarantee digital personhood are promising but require complex, unproven infrastructure and mass adoption. This remains an active area of research and development.
Nonetheless, until these new identity systems are perfected and widely adopted, governance systems must still defend against sybil attacks. One such system, already in use at Gitcoin, uses group membership signals to modulate funding payouts for Quadratic Voting (QV), which is a promising method for public funding of digital infrastructure. However, since QV awards matching funds, it is susceptible to coordinated sybil attacks where a user may create multiple identities to extract more funds than appropriate. Specifically, if individuals 1 through n contribute $c_{1}$ through $c_{n}$ to a project, then:
@@
Clearly, QV is highly susceptible to sybil attacks because the amplification of identities produces a squared payout.
@@
Augmented QV like Gitcoin’s “Passport” appears to offer a unique solution to a particular problem, but the proposed mechanisms are only robust against sybils to the extent that they require attestation to group membership. Miller, Weyl, and Erichsen [34] attempt to find (and formally define) appropriate ways of discounting collusion, based on identifiable group membership, and they offer several sketches for potential defence mechanisms. Three different clustering methods, a pooled centrality measurement, and an ‘eigen match’ are proposed. However, as mentioned, the clustering methods are only effective to the extent that privacy can be reduced by exposing group membership (in theory, a user could construct a Gitcoin ‘Passport’ using only privacy-preserving alternative identity systems, bootstrapping a DID, but this requires social behaviour and mass adoption). The proposed ‘offset match’ is recognised as somewhat crude, since it simply pools users and discounts contributions “to scale correlation scores so that everyone is ‘cared about’ equally strongly.” Finally, the ‘eigen match’ extends the pooling mechanism by focusing on individual behaviours using an adjacency matrix to construct a set of eigenvectors (somewhat similar to calculating Principle Components). However, as I discuss below, due to the anonymous nature of blockchain systems, solutions that rely on edge connections are not able to address singletons, which we assume are common and prevalent. The deep learning method we propose could replace an ‘eigen match’ component and provide more sophisticated QV sybil defence.
@@
Despite these challenges, the integration of blockchain technology and decentralized governance has brought about a paradigm shift in how commons security could be addressed. Unlike traditional commons, which primarily focus on monitoring and surveillance to prevent illegal resource extraction, digital commons, particularly those managed by DAOs, face a different set of challenges. In response to these evolving security needs, our method does not necessitate structural changes to existing governance systems and requires no new identity system. Instead, it emphasizes flexible, community-specified security responses to effectively counteract sybil attacks.
@@
However, this flexible security response raises a crucial question: what distinguishes a spam voter from a sybil voter, and is the distinction significant for DAOs? Conceptually, this distinction requires further analysis (is a spam voter just a sybil from the perspective of token voting?), but from a methodological standpoint, the distinction is pressing. If we define a sybil as a cluster of high dimensional embeddings, this implies a single hidden entity and a method for clustering based on similar vector embeddings seems appropriate. Conversely, if sybils are equated with spam (unwanted or harmful voting), targeting anomalous node embeddings for blocking becomes more relevant. Our approach adopts the former, differentiating between spam and sybils but more broadly it suggests a defense-in-depth strategy for DAOs, where outer layers target spam (anomalous voters) and inner layers focus on sybil threats (clusters of similar voters).

### v6

@@ Abstract
This research examines the polycentric governance of digital assets in blockchain-based Decentralized Autonomous Organizations (DAOs). It offers a theoretical framework and addresses a critical challenge facing decentralized governance by developing a method to identify Sybils, or spurious identities. Sybils pose significant organizational sustainability threats to DAOs and other, commons-based online communities, and threat models are identified. The experimental method uses an autoencoder architecture and graph deep learning techniques to identify Sybil activity in a DAO governance dataset (snapshot.org). Specifically, a Graph Convolutional Neural Network (GCNN) learned voting behaviours and a fast vector clustering algorithm used high-dimensional embeddings to identify similar nodes in a graph. The results reveal that deep learning can effectively identify Sybils, reducing the voting graph by 2-5%. This research underscores the importance of Sybil resistance in DAOs, identifies challenges and opportunities for forensics and analysis of anonymous networks, and offers a novel perspective on decentralized governance, informing future policy, regulation, and governance practices.
@@ 1   Introduction
Decentralized Autonomous Organizations (DAOs) represent a critical intersection of social behaviour and emerging technologies. Originating from blockchain architectures, DAOs have transcended their initial role in cryptocurrency governance to become complex social systems that defy traditional classification. In an attempt to elucidate their underlying structure and functions, this research offers a theoretical framework for DAOs and presents an inductive method to identify Sybils, thereby tackling a substantial issue facing the decentralized governance of digital assets.
A theoretical exploration of endogenous and exogenous models drawn from research on Online Communities, coupled with large scale empirical data and analysis, shapes the foundation of the neoinstitutional theory offered here. Such a theory is necessary to understand the purposes of DAOs and how they operate. In a nod to Elinor Ostrom’s work, we posits that DAOs can effectively govern Digital Common Pool Resources (DCPRs) through polycentric governance, despite their idiosyncratic challenges, if and only if they effectively manage Sybils. Sybils are fraudulent identities used to manipulate voting within digital decentralized governance (similar to ‘sockpuppets’ or ‘non-player characters’ in the system; the term originates in the early 2000s as part of cyber history’s dramatis personae). While traditional governance frameworks readily address the ‘one person, one vote’ principle, the anonymity provided by blockchain technologies complicates the influence of Sybils in decentralized governance.
To address this governance challenge, we used graph deep learning techniques to identify Sybil activity on a popular DAO governance forum, snapshot.org. Specifically, our method involves a Graph Convolutional Neural Net (GCNN) with multiple engineered feature layers that learned vector node embeddings from a voting graph to construct a ‘similarity’ subgraph, where similar nodes are identified using a fast k-means clustering algorithm (Facebook AI Similarity Search, FAISS) and relabelled. The similarity graph is then reduced by combining predicted Sybils. While this method is transductive and sensitive to feature development and learning hyperparameters, experimental results suggest that deep learning techniques can meaningfully detect Sybils and the resulting subgraph of ‘true’ identities is approximately 2-5% reduced. Social scientists, policy makers, system designers, and security professionals should consider these results when studying, designing, and implementing online communities and ensure that vulnerable governance practices are made resilient to Sybil attacks.
@@ 2   Deciphering DAOs: Advancing Research in Digital Collective Action
The key unifying factor among all DAOs is their use of, or responsibility for, a decentralized ledger, or “blockchain.” Central to each DAO is a shared digital asset on a blockchain, often referred to as a “token.” Recognizing the pivotal role of “token economics” in these novel, community-led platforms, we find a compelling area of research: an exploration of how these digitally augmented human collectives manage shared assets. This exploration draws our attention towards the concept of a “Common Pool Resource” [42] and our suggested adaptation, a “Digital Common Pool Resource” (DCPR). The further study of DCPRs could offer valuable insights into the nature and operation of DAOs.
The importance of DAO research also goes beyond academic interest. It offers an opportunity to challenge and extend our understanding of collective action, digital resource management, and decentralized governance. Further, studying DAOs offers insights into ongoing, cutting-edge experiments attempting to create robust, long-enduring institutional forms beyond the control (and regulation) of nation states. These efforts offer an organizational buttress for new and existing cryptocurrencies (typically seeking to decentralize) and frustrate criminal investigations. More optimistically and progressively, through detailed data analysis we can better comprehend the intricacies of DAOs and tap into their potential to reshape our digital environments, perhaps even as a form of “plural public” [31].
@@ 3   Emergence and Evolution of DAO Studies
As the discourse within DAO Studies evolves, we are seeing a shift from insular, participant-derived narratives towards the integration of established socio-economic theories that offer a broader context. A pivotal theoretical framework used to analyze these phenomena has been Hirschman’s paradigm of “Exit, Voice, and Loyalty]] [30]. In this model, ‘Exit’ refers to the option of leaving an organization or community when its quality deteriorates or when disagreements arise, while ‘Voice’ involves expressing dissatisfaction or working towards change from within the organization. ‘Loyalty,’ on the other hand, reflects a commitment to the organization that might inhibit ‘Exit’ and promote ‘Voice.’
@@
Similarly, many scholars have addressed how trust is transformed by DAOs. Much literature on the role of trust in blockchain technology presupposes that its function is to lower transaction costs within organizations, a sentiment summarized by Williamson [60] with his famous theory of Transaction Cost Economics (TCE). Halaburda et al.’s [27] research extends this argument by exploring how blockchain, as a ‘trust-reducing’ technology, can alter the dynamics of transaction governance in organizations, including DAOs. This work explores the intricacies of how blockchain’s inherent features, such as decentralization, transparency, and immutability, contribute to reducing uncertainties and transaction costs that are traditionally managed through hierarchical organizational structures. Halaburda’s insights are particularly relevant in the context of DAOs, which are presupposed to operate on principles of decentralized governance and collective decision-making. By reducing the need for intermediaries and centralized control, blockchain technology enables DAOs to operate with a higher degree of trust among participants (see the ‘Threat Model’ section 5.2.1 for a description of the motivation), fostering a collaborative environment that is less susceptible to opportunistic behaviors, a key concern in TCE.
Moreover, Halaburda’s perspective aligns with Hardin’s theory of trust, emphasizing how trust in DAOs is not merely a function of self-interest or moral commitment but also heavily influenced by the technological framework that underpins these organizations. This viewpoint is further developed by Lemieux [33], who reconstructs Hardin’s theory [29] to incorporate the role of blockchain technology in shaping trust dynamics. A subtle tripartite model of trust emerges from Hardin’s theory, where trust is “grounded in the truster’s assessment of the intentions of the trusted with respect to some action,” which is, typically, based in self-interest, moral commitment, and other idiosyncratic factors. Others, such as Wright and De Filippi [61], have focused on the ways technology transforms trust, drawing a parallel to the substitutability, or lack thereof, of law by code.
@@
In practice, however, most DAOs form as part of a process where a traditional startup team has built code and close personal trust first, and then formally declare intentions to decentralize. This core team, often with external advisors, legal counsel, and early stage traditional investors, then establish an off-chain DAO for representational—not direct—democratic control. A rare few will transition to on-chain governance where the “community” makes decisions directly (which still need to be implemented by software developers who have commit access to open source codebases). At this point in development, decentralized governance becomes complex, particular, and institutional in nature. Arbitrum offers a case study of this developmental path, also known as “Progressive Decentalization” (see [17]). The institutionalization of Arbitrum, for example, involved the launch of an off-chain DAO, a governance token launch and “airdrop,” a quasi-legal “constitution,” and a progressive decentralization of codebases led by an independent and legal foundation.
@@ 3.1   Quantitative and Empirical DAO Research
For comparison, we performed HHI (see Figure 1), Gini (see Figure 2) and $n$-top players (see Figure 3) analyses on our dataset of 7.2m votes.
@@
In terms of governance, in the Snapshot dataset Bellavitis, Fisch, and Momtaz found that 15% of DAOs have a token ownership threshold to make governance proposals, 160 votes on average per voter, and the proposal approval rate has hovered around 90%. However, the voter counts are naive and do not discount the considerable noise in DAO governance nor the obvious role of Sybils, so while the results offer a helpful baseline, we question the accuracy of the findings. For instance, using slightly less naive counting methods on a much larger Snapshot dataset of 59,939 proposals, we find approximately 100 average votes per voter.
@@ 3.2   Existing Research Challenges for DAO Studies
The theoretical underpinnings of DAO studies are relatively undeveloped, with the initial body of literature often drawing from a ‘native’ theory of technological utopianism that is self-referential and uncritical. This self-referential use of theory may lead to an echo chamber effect that focuses attention on decentralization and autonomy at the expense of richer social phenomena.
@@ 4   Online Communities
4.2   Towards Polycentric Governance
@@
4.2.1 Digital Common Pool Resources
Ostrom’s ideas about polycentric governance can be adopted and adapted to digital assets and their unique governance challenges. Here, we present the novel idea of “Digital Common Pool Resources” (DCPRs) as a digital adaptation of her theory of Common Pool Resources (CPRs). In this context, resources are not physical but digital, like codebases, online communities, or cryptographic assets. Ostrom’s principles find renewed relevance in the rapidly evolving world of blockchain and cryptocurrency, where decentralized, community-based control is paramount. Scholars such as De Filippi and Hassan [24] have further explored how blockchain technologies align with Ostrom’s principles, offering new opportunities for cooperative governance. Ostrom’s neoinstitutional approach provides a foundational framework for understanding self-governance in decentralized digital systems, focusing on the human-centered aspects of these technologies. This approach emphasizes collective action, reciprocity, and shared governance, aligning closely with the core philosophies of decentralization in the digital age.
@@ 4.2.1 Digital Common Pool Resources
By applying these principles to the decentralized governance of Web3, the creation and management of digital commons can mirror the successful self-governance of tangible resources. Our approach assumes that DAOs are responsible for DCPRs and we contribute a method for identifying Sybils to mitigate a key security challenge facing digital assets, helping ensure that these commons can be managed effectively and sustainably.
@@ 5.1   Ethical considerations
It should be noted that using deep learning methods to analyze anonymous users on public financial and data records poses significant ethical risks and such methods should be considered a form of social network deanonymization. Even though data are ostensibly public, individuals may not have consented or even be aware that their financial transactions could be linked back to them in an identifiable way. This becomes particularly troubling given that financial data can reveal sensitive aspects of an individual’s spending habits, income sources, and social networks—that they may not want publicly disclosed. In light of previous findings [16] on ethical research risks for cryptocurrencies, even when lacking real “human subjects,” research without consent exacerbates ethical concerns. Moreover, researchers undertaking this type of analysis must be acutely aware that their methods could be weaponized to infringe upon privacy or even perpetrate crimes.
@@ 5.2   Sybil Resistance and Defence in Depth for Community Design
These efforts, largely emerging from computer science, parallel the evolution of Ostrom-style commons management practices emerging in DAOs. DAOs continue to experiment with alternative identity systems, including Proof of Personhood (PoP), Proof of Participation (PoP), Decentralized IDs (DIDs), and Self Sovereign IDs (SSIDs). New identity systems that guarantee digital personhood are promising but require complex, unproven infrastructure and mass adoption. This remains an active area of research and development.
5.2.1 Threat Model
Implicitly, we have assumed that Sybils threaten governance because they can be used to mask influence, but we have yet to define the specific threat model. Our threat modeling must accept that anonymous or pseudonymous voting is seen as essential in these communities. Thus, without an identity infrastructure, an individual can easily create multiple identities to have their vote counted more than once. Since votes are tokens, an individual can create multiple identities by simply splitting their coins among multiple wallets. Given widespread economic inequality (the ‘whales’ identified in [26]), ‘buying’ a proposal is a realistic threat against DAOs (see Figure 5). This attack can be multiplied by buying tokens on margin, voting, and returning the borrowed tokens—a low-cost flash attack. This overall arrangement—no durable identities and unrestricted coin voting—threatens the long-term sustainability of DAOs, which can be drained of resources, compelled to make destructive strategic decisions, or hijacked.
Beyond plutocratic voting and community and resource misalignment, DAOs face specific threats associated with Sybils. For example, many DAOs have embraced Quadratic Voting (QV) as a way to more equitably distribute influence. Similarly, public-funding websites like Gitcoin have also embraced QV to modulate funding payouts. Specifically, if individuals 1 through n contribute $c_{1}$ through $c_{n}$ to a project, then:
@@ 5.2.1 Threat Model
However, since QV awards matching funds or voting power, it is susceptible to coordinated Sybil attacks where a user may create multiple identities to extract funds or influence governance outcomes. QV is susceptible to Sybil attacks because the amplification of identities produces a squared payout, which can be used to drain resources or exert additional governance influence.
@@
Augmented QV like Gitcoin’s “Passport” appears to offer a unique solution to a particular problem, but the proposed mechanisms are only robust against Sybils to the extent that they require attestation to group membership. Miller, Weyl, and Erichsen [38] attempt to find (and formally define) appropriate ways of discounting collusion, based on identifiable group membership, and they offer several sketches for potential defence mechanisms. Three different clustering methods, a pooled centrality measurement, and an ‘eigen match’ are proposed. However, as mentioned, the clustering methods are only effective to the extent that privacy can be reduced by exposing group membership (in theory, a user could construct a Gitcoin ‘Passport’ using only privacy-preserving alternative identity systems, bootstrapping a DID, but this requires social behaviour and mass adoption). The proposed ‘offset match’ is recognised as somewhat crude, since it simply pools users and discounts contributions “to scale correlation scores so that everyone is ‘cared about’ equally strongly.” Finally, the ‘eigen match’ extends the pooling mechanism by focusing on individual behaviours using an adjacency matrix to construct a set of eigenvectors. However, as I discuss below, due to the anonymous nature of blockchain systems, solutions that rely solely on edge connections are not able to address singletons, which we assume are common and prevalent. The deep learning method we propose could replace an ‘eigen match’ component and provide more sophisticated QV Sybil defence.
Figure 4: Unique Voting ‘Strategies’ used by 100 DAOs.
On Snapshot.org, DAOs use a complex assemblages of bespoke voting “strategies” to ultimately calculate a user’s voting power (on-chain voting like MakerDAO uses multiple smart contracts to determine voting power). For example, 81 of 100 DAOs use a simple ERC-20 token ‘balance’ strategy (see Figure 4). However, other strategies can include token delegation and arbitrary Solidity chain code execution to calculate voting power. As such, it is non-trivial to calculate the cost to ‘buy’ a governance proposal, but with a few assumptions we can determine some baseline minimum costs. Specifically, we used the Covalent API to calculate the historical price of each ERC-20 voting token (assuming each user rationally maximized their voting power, which we know to be empirically false), and summed the costs to purchase ten random governance proposals for three example DAOs (see Figure 5).
Figure 5: Minimum cost to purchase governance proposals in three example DAOs.
The typical free/libre and open source (FLOSS) alternative to decentralized governance (as with DAOs) has distinct challenges when tasked with sustainably and securely managing digital commons. The traditional FLOSS governance models (Linux’s benevolent dictator Linus, Wikipedia’s altruistic editors, Apache’s Foundation model, etc.) all suffer from a reliance on personal trust or legal structuring, neither of which are robust against malicious strategic attack. For the sustainability of digital commons, the decentralization of human decision making is essential, which is why we expect to see more DAOs in the future.
Monero, the leading crime and privacy cryptocurrency, offers an interesting case study of the challenges facing traditional FLOSS management of valuable digital resources. Like many traditional FLOSS software projects, Monero relies on a trusted core team of developers with commit access to a code repository. However, despite years of successful development and increasing use, Monero has recently encountered a potentially fatal organizational issue; a theft of donations (practically all cryptocurrencies encounter such threats eventually). In late 2023 a Monero wallet maintained by several core project members was drained. Accusations of internal theft pointed to a core member (@Luigi) who had recently been arrested (apparently on unrelated charges). This set of events shook up the Monero project which led to a transition to a decentralized governance model. The long term success of Monero hinges on this transition.
In an effort to better understand the threat model facing traditionally-organized (FLOSS) cryptocurrencies like Monero, we used the GitHub API to collect all (2,984) of Monero’s ‘Issues’ and the associated comments. Each GitHub issue with its comments was then processed by a LLM (GPT-4 Turbo), prompted to identify the main issues, relevant actors, sentiment, and to categorize as either governance, security, technical, or financial. The result was further analysed by concatenating the summaries and passing these back into the LLM for further summarizing (to accommodate input token restrictions for the LLM, we vectorized the summaries, trained embeddings with a BERT pre-trained model, and clustered into eight, appropriately sized chunks of input text). Some notable threat issues emerged from this bootstrapped history (see Appendix Appendix: Retrieval Augmented Generation (RAG) of Software Engineering Issues for Monero Cryptocurrency: Hierarchical GPT-4 Turbo for Automated Histories*), including debates about infrastructure changes, community engagement and moderation, and finally, two security issues that reflect broader issues of transparency and governance. This hierarchical Retrieval Augmented Generation (RAG) method can be used at scale for rapidly generating governance histories.
@@
Despite facing myriad governance challenges (many of which are still unknown), DAOs bring about a paradigm shift in how commons can be managed. Unlike traditional FLOSS, DAOs significantly alter expectations of transparency and trust, but demand stringent security practices. Like traditional commons, DAOs require monitoring and surveillance to prevent illegal resource extraction. But, unlike traditional commons, DAOs face a different set of challenges, especially Sybils. In response to these evolving security needs, our method of Sybil identification does not necessitate structural changes to existing governance systems and requires no new identity system. Instead, it emphasizes flexible, community-specified security responses to effectively counteract Sybil attacks.
@@
However, this flexible security response raises a crucial question: what distinguishes a spam voter from a Sybil voter, and is the distinction significant for DAOs? Conceptually, this distinction requires further analysis (is a spam voter just a Sybil from the perspective of token voting?), but from a methodological standpoint, the distinction is pressing. If we define a Sybil as a cluster of high dimensional embeddings, this implies a single hidden entity and a method for clustering based on similar vector embeddings seems appropriate. Conversely, if Sybils are equated with spam (unwanted or harmful voting), targeting anomalous node embeddings (noise) for blocking becomes more relevant. Our approach adopts the former, differentiating between spam and Sybils but more broadly it suggests a defense-in-depth strategy for DAOs, where outer layers target spam (anomalous voters) and inner layers focus on Sybil threats (clusters of similar voters).